Friday, 21 August

Friday, 21 August

Invited Lectures

\underline{14:00 – 14:45}

04 Mikhail M. Kapranov, Northwestern University, Evanston, USA Operads and algebraic geometry (I: p.65) Chair: Enrico Arbarello

MA 005

05 Simon K. Donaldson, Stanford University, Stanford, USA Positive line bundles in symplectic geometry (I: p.69) Chair: Ulrich Pinkall

MA 001

06 Alexander N. Dranishnikov, University of Florida, Gainesville, USA and Steklov Math. Institute, Moscow, Russia Dimension theory and large Riemannian manifolds (I: p.79) Chair: Alexander S. Mishchenko

MA 004

09 Shuhei Hayashi, Waseda University, Tokyo, Japan Hyperbolicity, stability and the creation of homoclinic points (I: p.113) Chair: Jack Carr

H 1012

11 Fedor Smirnov, CNRS, LPTHE, Université Pierre et Marie Curie, Paris, France Space of local fields in integrable field theory and deformed Abelian integrals (I: p.137) Chair: Andreas Fring

H 1028

14 Emo Welzl, ETH Zürich, Zürich, Switzerland Halving Point Sets (I: p.175) Chair: Helmut Alt

H 2032

\underline{15:00 – 15:45}

04 Victor Vadimovich Batyrev, Universität Tübingen, Tübingen, Germany Mirror symmetry and toric geometry (I: p.60) Chair: Yujiro Kawamata

MA 005

05 François Labourie, Université Paris-Sud, Orsay, France Large Group Actions on Manifolds (I: p.73) Chair: Ulrich Pinkall

MA 001

06 Michael Hartley Freedman, Microsoft Corporation, Redmond, USA Topological views of computational complexity (I: p.80) Chair: Robert D. Edwards

MA 004

09 Michael Herman, CNRS, Université Denis Diderot / Paris 7, Paris, France Some open problems in dynamical systems (I: p.114) Chair: Thomas Bartsch

H 1012

11 Roberto Schonmann, UCLA, Los Angeles, USA Metastability and the Ising model (I: p.136) Chair: Christian Borgs

H 1028

14 Miklós Ajtai, IBM Almaden Research Center, San Jose, USA Worst-case Complexity, Average-case Complexity, and Lattice Problems (I: p.165) Chair: Helmut Alt

H 2032

\underline{16:15 – 17:00}

02 William M. Kantor, University of Oregon, Eugene, USA Simple Groups in Computational Group Theory (I: p.41) Chair: Charles C. Sims MA 005 07 Vera Serganova, University of California, Berkeley, USA On representations of Lie superalgebras (I: p.95) Chair: Olivier Mathieu

MA 004

08 Michael Christ, University of California, Berkeley, USA Singularity and Regularity — Local and Global (I: p.97) Chair: Daryl Geller

MA 001

10 Gustavo Alberto Ponce, University of California, Santa Barbara, USA On nonlinear dispersive equations (I: p.125) Chair: Mikhail Shubin

H 1012

12 Mark Freidlin, University of Maryland, College Park, USA Random and Deterministic Perturbations of Nonlinear Oscillators (I: p.141) Chair: Jürgen Gärtner

H 1028

16 & 17 William J. Cook, Rice University, Houston, USA The Traveling Salesman Problem (I: p.186) Chair: Günter M. Ziegler

H 2032

\underline{17:15 – 18:00}

02 Aleksandr Pukhlikov, Institute for Systems Analysis, Moscow, Russia Birational automorphisms of algebraic varieties (I: p.42) Chair: Victor Vadimovich Batyrev

MA 005

07 Shahar Mozes, Hebrew University, Jerusalem, Israel Products of trees, lattices and simple groups (I: p.93) Chair: Peter Sarnak

MA 004

08 Hart F. Smith, University of Washington, Seattle, USA Wave Equations with Low Regularity Coefficients (I: p.107) Chair: Gerald B. Folland

MA 001

10 Frédéric Hélein, CMLA, ENS-Cachan, Cachan, France Phenomena of compensation and estimates for partial differential equations (I: p.121) Chair: Gunther A. Uhlmann

H 1012

12 Jean-François Le Gall, DMI — École Normale Supérieure, Paris, France Branching processes, random trees and superprocesses (I: p.147) Chair: Donald Dawson

H 1028

17 Jan C. Willems, University of Groningen, Groningen, The Netherlands Open Dynamical Systems and their Control (I: p.195) Chair: Fritz Colonius H 2032